016 (part 1 of 2) 10.0 points A neutron in a reactor makes an elastic headon collision with the nucleus of an atom initially at rest. Assume: The mass of the atomic nucleus is about 13.7 the mass of the neutron. What fraction of the neutron’s kinetic energy is transferred to the atomic nucleus? 017 (part 2 of 2) 10.0 points If the initial kinetic energy of the neutron is 8.06 × 10−13 J, find its final kinetic energy. Answer in units of J.
Answer #1

a)

m = mass of neutron

M = mass of nucleus = 13.7 m

vi = initial velocity of neutron before collision = ?

Vi = initial velocity of nucleus before collision = 0 m/s

vf = final velocity of neutron after collision

Vf = final velocity of nucleus after collision

using conservation of momentum

m vi + M Vi = m vf + M Vf

(m) vi + (13.7 m) (0) = m vf + (13.7 m) Vf

vi = vf + (13.7) Vf     

vi - vf = (13.7) Vf                                                  eq-1

using conservation of kinetic energy

m v2i + M V2i = m v2f + M V2f

(m) v2i + (13.7 m) (0)2 = m v2f + (13.7 m) V2f

v2i = v2f + (13.7) V2f

v2i - v2f = (13.7) V2f

( vi - vf )( vi + vf ) = (13.7) V2f

(13.7) Vf  ( vi + vf ) = (13.7) V2f

vi + vf = Vf                                             eq-2

adding eq-1 and eq-2

( vi - vf ) + ( vi - vf ) = (13.7) Vf + Vf

2 vi = 14.7 Vf

vi = 7.35 Vf

fraction of energy transferred is given as

fraction = (0.5) M V2f /((0.5) m v2i )

fraction = M V2f /(m v2i )

fraction = (13.7 m) V2f /(m (7.35 Vf )2)

fraction = 13.7/(7.35)2 = 0.25

b)

KEni = initial kinetic energy of neutron = 8.06 x 10-13 J

KEnf = initial kinetic energy of neutron

KEN = final kinetic energy of nucleus

Given that

fraction = KEN /KEni

0.25 = KEN /(8.06 x 10-13)

KEN = 2.02 x 10-13 J

also

KEni = KEnf + KEN

8.06 x 10-13 = KEnf + 2.02 x 10-13

KEnf = 6.04 x 10-13 J